if M (x,y) = (x+y, x+y ) 

Find the matrix representation of M with respect to (v1,v2) where v1= (1,1) v2 = (1,3) 


1
Expert's answer
2021-06-08T12:03:58-0400

Consider a vector "l=xv_1+yv_2=(x+y,x+3y)". We act on a vector with a matrix "M" and receive: "(2x+4y,2x+4y)=(2x+4y)v_1". Thus, the matrix "M" with respect to "(v_1,v_2)" has a representation: "M_{v_1,v_2}=\left(\begin{array}{cc}2&4\\0&0\end{array}\right)".